What is the molarity of a solution that contains 0.350 g of barium fluoride, BaF2, in 0.50 L of solution
1
Expert's answer
2020-12-01T09:37:05-0500

Mass of BaF2= 0.350g

Molar mass of BaF2= 137.34 + (19x2)= 175.34gmol-1

Moles of BaF2= mass/molar mass

Moles of BaF2= 0.350/175.34= 2.0 x 10-3mol

Volume of solution = 0.50L

Molarity = moles of solute/litres of solution

Molarity = 2.0 x10-3mol/ 0.5L

Molarity= 4.0 x 10-3M